### v1

Abstract
Foundation Models (FMs) are large-scale, pre-trained AI systems that have revolutionized natural language processing and computer vision, and are now advancing geospatial analysis and Earth Observation (EO). They promise improved generalization across tasks, scalability, and efficient adaptation with minimal labeled data. However, despite the rapid proliferation of geospatial FMs, their real-world utility and alignment with global sustainability goals remain underexplored. We introduce SustainFM, a comprehensive benchmarking framework grounded in the 17 Sustainable Development Goals with extremely diverse tasks ranging from asset wealth prediction to environmental hazard detection. This study provides a rigorous, interdisciplinary assessment of geospatial FMs and offers critical insights into their role in attaining sustainability goals. Our findings show: (1) While not universally superior, FMs often outperform traditional approaches across diverse tasks and datasets. (2) Evaluating FMs should go beyond accuracy to include transferability, generalization, and energy efficiency as key criteria for their responsible use. (3) FMs enable scalable, SDG-grounded solutions, offering broad utility for tackling complex sustainability challenges. Critically, we advocate for a paradigm shift from model-centric development to impact-driven deployment, and emphasize metrics such as energy efficiency, robustness to domain shifts, and ethical considerations.
keywords:
Foundation Models, Earth Observation, Sustainable Development Goals, Geospatial AI, Benchmarking
Introduction
The Sustainable Development Goals (SDGs) are a collection of 17 global goals set by the United Nations General Assembly in 2015. They aim to achieve a better and more sustainable future by addressing global challenges related to poverty, inequality, climate change, environmental degradation, peace, and justice. The SDGs offer a shared framework for countries, organizations, and individuals to advance an economically, socially, and environmentally sustainable world by 2030. Earth observation (EO) techniques significantly contribute to sustainable development by providing critical data and insights that support the achievement of SDGs [1, 2, 3]. EO contributes to global challenges related to poverty, inequality, climate change, environmental degradation, peace, and justice. From monitoring climate change, biodiversity, and water resources to enhancing food security, urban planning, and disaster resilience [4, 5, 6, 7], EO provides critical information for tracking environmental and socio-economic trends. Through supporting evidence-based policy and providing an enormous amount of up-to-date data, EO accelerates progress toward a more sustainable, equitable, and resilient future.
Exploiting EO data at scale requires automated tools; foundation models (FMs) offer a powerful solution for handling the diverse, multimodal spatio-temporal data typical of geospatial applications. FMs are large-scale artificial intelligence (AI) models that are pre-trained on extensive, diverse, and often multimodal datasets using self-supervised learning, which involves learning from raw, unlabeled data via proxy tasks, or weakly supervised learning, which utilizes noisy or incomplete labels [8]. FMs contain billions of parameters (e.g., Google’s SegGPT with over 3 billion and Meta’s DINOv2 with around 1 billion parameters), and leverage vast computational resources to learn detailed patterns and representations from massive datasets, which enables them to generalize across a wide range of downstream tasks. With this definition, they serve as a universal backbone to support a wide range of applications across domains such as natural language processing (NLP), computer vision, and (geo)scientific computing. The term foundation emphasizes their role in providing a pre-trained, transferable representation that can be fine-tuned or adapted for domain-specific applications. Geospatial FMs usually contain hundreds of millions of parameters and are pre-trained on vast airborne and spaceborne data. They can then be adapted to specific downstream tasks, e.g., land cover classification, change detection, or disaster response mapping, without requiring separate models to be trained from scratch for each task [9]. Key characteristics of FMs include: (1) Scalability through efficient, distributed training; (2) Generalization via transferable representations across tasks; and (3) Few-/Zero-shot Learning, enabling adaptation with little or no labeled data.
Figure 1: Summary of the SustainFM Benchmark. By bringing together domain experts and AI specialists (SDG 17: Partnerships for the Goals), and providing a collection of 16 datasets aligned with the first 16 SDGs, from six continents, spatial resolutions ranging from 0.5 m to 30 m, and multiple tasks grounded in real-world applications, SustainFM aims to provide a testbed to analyze the applicability and real impact of foundation models.
Since 2023, the EO community has witnessed an explosion in the development of FMs, evolving from single-modal [10, 11, 12] to multi-modal [13, 14, 15, 16] architectures that integrate diverse data sources such as optical imagery, SAR, and auxiliary geospatial data. These advancements aim to enable more comprehensive and robust downstream applications. However, the rapid growth in the number of FMs over such a short period has led to a saturation of new architectures, resembling the clear trajectory we witnessed in the EO community during the deep learning boom around 2016, often marked by incremental improvements and limited consideration of their societal, environmental, and economical impact. Despite advancements in computational efficiency and scalability, a fundamental gap remains in grounding model development with global sustainability objectives. To see the actual impact of FMs on real-world challenges, such as climate monitoring, resource management, and humanitarian applications, there needs to be a stronger focus on the deployment phase (downstream tasks), which is widely regarded as the ultimate goal of Responsible AI in EO [17].
To address this, we must shift the emphasis from mere model development to the deployment phase, where FMs are evaluated based on their impact on complex socio-environmental contexts. Performance evaluation should go beyond traditional benchmarks and include real-world scenarios, social good, and ethical considerations in the deployment stage. Furthermore, the interdisciplinary views of both AI and domain experts are required to ensure that FMs are not only optimized for accuracy but also for their contribution to long-term societal benefits. Without this shift, we risk an unsustainable model development cycle with no real-world impacts.
In this paper, we move beyond the current hype of FMs to adopt an evidence-based approach to answer this question:
1. Do we need FMs, or are traditional, purpose-specific deep learning models sufficient for EO tasks?
2. What evaluation criteria best capture the effectiveness and responsibility of FMs in EO, beyond task-specific accuracy?
3. Can FMs be systematically leveraged to support scalable, SDG-grounded solutions for geospatial challenges?
Through systematic evaluation across key criteria, including transferability, data efficiency, energy costs, and operational impact, we aim to provide a grounded, rigorous assessment that challenges assumptions and clarifies the actual added value, if any, of FMs in this domain. Ultimately, our goal is to contribute to global sustainability objectives and promote social good, a key aspect that has been undermined in many current geospatial FMs. To evaluate FMs in EO rigorously, we introduce SustainFM, a benchmark aligned with all 17 SDGs that provides a new dataset, leverages pretrained FMs, and evaluates them for global sustainability, as conveyed in Fig. 1. This dataset spans a wide range of geographies (over 200 countries and regions across continents, as shown in global distribution in Fig. 2), sensor types, and spatial resolutions (from 0.5 m to 30 m) to reflect the diversity and complexity of global sustainability challenges. SustainFM combines 16 real-world tasks (dedicated to the first 16 SDGs), including asset wealth prediction, child and women’s health estimation, solar farm and settlement mapping, and detecting industrial smoke, marine pollution, and conflict-related damage. These tasks cover key learning tasks such as regression, change detection, segmentation, and classification, and are tackled using a variety of sensors such as Landsat, Sentinel-1/2, VIIRS, Gaofen-2, Harmonized Landsat and Sentinel-2 (HLS), Google Earth, and PlanetScope. By encompassing domains like climate resilience, food security, infrastructure monitoring, and humanitarian response, SustainFM enables a comprehensive assessment of FMs across transferability, data efficiency, and real-world utility. By bringing together experts with diverse backgrounds, from domain scientists and remote sensing specialists to AI ethicists, machine learning researchers, and statisticians, we contribute to SDG-17 (Partnerships for the Goals) and offer critical yet comprehensive insights into whether FMs can contribute to global sustainability, moving beyond academic benchmarks to real-world impact.
Figure 2: Geographic coverage of the SustainFM benchmark, which spans over 200 regions across 6 continents to support a wide range of EO tasks aligned with the SDGs.
Empirical Evidences
We compare various geospatial FMs and traditional deep models using the results obtained on the SustainFM Benchmark. SustainFM includes data from six continents and incorporates data from a wide range of satellite sensors with spatial resolutions from 0.5 to 30 meters. To ensure task diversity, we include a variety of AI tasks such as regression (Reg.), semantic segmentation (Seg.), change detection (CD), and classification (Cls.), all of which are grounded in real-world applications, and includes a total of 723,555 image patches with patch sizes ranging from 120×120 to 1024×1024 to reflect reality. We provide detailed information about the building blocks of SustainFM, its diversity, and investigated models in Section 5, Methodology.
Table 1 presents the performance of various FMs (i.e., CROMA [18], RemoteCLIP [19], SoftCon [20], DOFA [21], GFM-Swin [22], Prithvi [23], SpectralGPT [10], ScaleMAE [24], SSL4EO-S12 [25], and SatlasNet [26]) and traditional deep models (i.e., Vision Transformer (ViT) [27] and Resnet-50 [28]) across the SustainFM tasks. We report performance using root mean square error (RMSE) for regression and mean F1 score (mF1) for classification, to capture both the prediction of continuous values and the identification of categorical classes.
Table 1: Performance of Foundation Models and Traditional Methods across SustainFM tasks. Metrics are RMSE (↓) or mF1 (↑) as appropriate. The best performance per column is shown in bold, and the second-best is underlined.
Model #1 #2 #3 #4 #5 #6 #7 #8 #9 #10 #11 #12 #13 #14 #15 #16
CROMA 1.39 84.10 35.01 2.75 2.10 84.94 96.52 69.78 93.10 0.99 91.92 73.87 89.66 70.73 87.50 96.71
[…]
GFM-Swin 1.45 87.74 34.74 2.71 2.08 87.89 94.86 75.83 93.59 1.04 93.78 78.05 85.05 74.14 88.62 97.93
[…]
RemoteCLIP 1.41 86.23 35.03 2.72 2.15 87.06 94.51 74.14 95.02 1.07 92.74 79.32 85.10 71.18 87.98 97.31
[…]
Scale-MAE 1.48 86.56 34.77 2.75 2.16 88.34 95.52 69.08 94.68 1.10 94.94 77.21 86.11 68.18 89.69 96.40
[…]
SSL4EO-S12 1.51 82.29 34.90 2.89 2.34 83.30 95.87 69.16 93.65 0.96 91.45 70.28 88.89 64.73 87.12 95.62
[…]
ViT 1.46 77.82 34.74 2.86 2.20 74.31 90.58 59.93 93.17 1.10 92.73 74.52 79.79 34.67 82.71 93.17
[…]
Despite the diversity in pretraining strategies, most FMs utilize transformer-based architectures, with a notable consideration of ViT [27]. Therefore, we adopt ViT as a baseline architecture to establish a performance reference. We train ViT from scratch on each subset of the SustainFM benchmark, treating it as a conventional deep learning model. Its performance is then compared against pre-trained and fine-tuned FMs. In addition, we include a traditional deep learning model, ResNet [28], as a representative of convolutional neural networks (CNNs), widely used for geospatial tasks over the past decade, as the second traditional model.
We freeze the FM encoder weights and fine-tune only the decoder parameters [29]. FMs, with their pretrained encoders frozen and decoders fine-tuned on SDG-specific tasks, generally exhibit superior performance across diverse tasks than trained-from-scratch traditional models. This advantage comes from their ability to leverage large-scale pretraining, allowing them to adapt and generalize more effectively across various SDG challenges. Additionally, based on our experiments, FMs converge faster in the training stage than traditional models, as their pretraining will enable them to start with established knowledge. This advantage helps to reduce the time and data needed for fine-tuning on SDG tasks. On the other hand, traditional models like ViT and ResNet-50, which are trained from scratch for each SDG task, may require more extensive tuning to achieve optimal performance, especially for broader generalization. While FMs often excel in flexibility, faster convergence, and achieving state-of-the-art results, traditional models can still perform competitively on a few tasks.
To assess different models’ environmental impact and sustainability, we employ the CodeCarbon library [30] to monitor energy consumption during training and inference, and show the results in Table 2. Specifically, we compare the energy footprint of decoder-only fine-tuning with full model fine-tuning, alongside the energy demands of baseline and traditional architectures (ViT and ResNet-50) trained from scratch. All models are evaluated on the OMBRIA dataset (Life on Land, SDG-15). In general, fine-tuning only the decoder of an FM can slightly boost performance while using less energy and producing lower CO2 emissions than fine-tuning the entire model. This is because it requires fewer gradient computations and backpropagation steps. This relative difference varies substantially across all models, ranging from +40% to as much as +168%, as can be seen in the last column of Table 2, highlighting the considerable environmental benefits of updating only the decoder. For context, full fine-tuning of SpectralGPT produces around 270g of CO2 emissions. This is roughly equivalent to driving a car for 0.7 miles, or less than 1% of the average American’s daily per-person emissions (39 kg/day) [31]. Complete training for traditional deep learning models usually results in higher energy consumption and CO2 emissions compared to partially fine-tuned FMs, since conventional models require training from scratch on large datasets over many iterations. In contrast, FMs benefit from pretraining on diverse data; They can be adapted to downstream tasks using a small amount of labeled data and fewer updates, which makes them more energy-efficient and scalable in the long run.
Table 2: Comparison of Finetuning the FM decoder only and Finetuning the Full Parameters. ViT and ResNet-50, the representatives of traditional deep learning models, are trained from scratch for SDG 15 (i.e., the OMBRIA dataset for flood mapping).
Model Finetuning Decoder Only Finetuning Full Parameters
mF1 $\uparrow$ Training Energy (kWh) Inference Energy (kWh) CO2 Emission (kg) mF1 $\uparrow$ Training Energy (kWh) Inference Energy (kWh) CO2 Emission (kg) CO2 Emission Relative Difference (%)
[…]
DOFA 89.328 0.0359 0.00029 0.02121 89.262 0.05288 0.00029 0.03111 +39.46
[…]
Prithvi 87.864 0.03473 0.00029 0.02104 87.964 0.05246 0.00028 0.03071 +61.11
[…]
SatlasNet 87.416 0.03428 0.00041 0.02028 89.19 0.09296 0.0004 0.05436 +168.05
[…]
SpectralGPT 87.663 0.19017 0.00142 0.11206 87.75 0.45666 0.0016 0.26864 +139.73
[…]
SoftCon 88.191 0.0417 0.00033 0.02498 88.089 0.06883 0.00041 0.04055 +62.33
Trained from Scratch for traditional deep models
[…]
ResNet-50 - - - - 82.756 0.03782 0.00011 0.02346 -
Discussions
Table 3 summarizes the strengths and weaknesses of FMs (decoder fine-tuned) compared to traditional models in terms of performance, efficiency, and sustainability criteria. Here, we discuss our key findings, as well as principles and considerations for geospatial FMs.
Table 3: Comparison between Foundation Models (decoder fine-tuned) and Traditional Deep Models (trained from scratch), across performance, efficiency, and sustainability metrics. Color-coded markers indicate relative performance: ✓best, ✓competitive, and ✗ limitation.
Assessment Criteria FMs (Decoder Fine-tuned) Traditional Models
[…]
Training Efficiency (kWh) ✓lower training cost ✓moderate
[…]
CO2 Emissions (kg) ✓generally low ✓moderate
[…]
Generalization across tasks ✓broad task adaptability ✗ narrower scope
[…]
Key Findings
FMs deliver strong performance, but are not always superior to traditional approaches. Our results indicate that it is impossible to make a universal claim about the superiority of FMs over traditional deep learning models (e.g., ViT) in terms of accuracy. Performance varies depending on the dataset, task, and fine-tuning strategy. Nonetheless, in many of our experiments, FMs outperformed their traditional counterparts.
Transferability offers a step toward energy-efficient AI across domains. A commonly held belief is that FMs are inherently resource-intensive. However, our results suggest a more balanced picture. Once pretrained, FMs can be fine-tuned with minimal labeled data, which offers an energy-efficient alternative to training models from scratch. This transferability makes them especially attractive for zero-shot or few-shot learning scenarios, where resource constraints or data scarcity are significant limitations.
FMs enable scalable solutions for global sustainability. One of the most promising aspects of FMs is their potential to serve as general-purpose tools across geospatial applications. Their ability to generalize across regions, tasks, and modalities makes them particularly suited to addressing complex, large-scale problems related to the SDGs. This cross-domain flexibility allows broader, interdisciplinary adoption among scientists and practitioners.
High generalization and fast convergence make FMs a scalable alternative to traditional models. Traditional deep learning models typically require task-specific training from scratch to achieve competitive performance, which results in considerable computational and energy costs for each new use case. In contrast, FMs exhibit strong generalization capabilities across various tasks, often requiring only minimal fine-tuning. This enables faster convergence and lower overall resource consumption in the long term.
Beyond the Hype
Aligning models with the nature of EO: Training FMs from scratch requires an enormous amount of time and computational resources, which make it largely impractical for benchmarking studies like ours. Given these constraints, we focused on fine-tuning existing pre-trained models. Accordingly, the energy use reported in our work reflects only this fine-tuning phase. We strongly encourage researchers to clearly report CO2 emissions or other environmental impact measures whenever they introduce new FMs to provide the whole picture. This transparency is essential to support responsible AI development and help the community move toward more sustainable practices in training and using large models.
The development of geospatial FMs demands a methodological shift that reflects the unique nature of EO data. As argued in [32], EO data is not just another visual modality; They include unique properties, such as spectral and temporal richness, geolocation, physical constraints, and sensor-specific noise patterns, which are not adequately addressed by conventional vision-based pretraining objectives such as masked image modeling or contrastive learning or the models themselves [33]. As a result, downstream geospatial tasks such as change detection, time-series forecasting, and biophysical retrieval may suffer from a misalignment between model pretraining and real-world task demands. Moreover, relying on standard computer vision architectures, such as UperNet or Siamese UperNet, can hide the real impact of task-specific reasoning in geospatial applications. To move the field forward, we advocate for models that capture the specific nature of geospatial data. This includes using physics-informed pretraining that brings in domain knowledge, learning methods that combine different types of data (like spatial, spectral, and temporal), and flexible decoder designs with suitable loss functions that match the needs of specific geospatial tasks.
Furthermore, the importance of scale must not be overlooked. Though originally developed for language models, scaling laws emphasize the importance of maintaining an appropriate data-to-parameter ratio, typically ranging from 5:1 to 20:1, to achieve generalizable performance [34]. As highlighted in [33], many geospatial FMs are undertrained relative to their size, limiting their generalization capacity despite architectural scale. Smaller models risk underfitting complex Earth patterns, while larger models require correspondingly large and diverse training datasets to fulfill their potential.
Responsible practices with building fair and sustainable models: The environmental and ethical responsibilities tied to geospatial FM development are substantial. The carbon footprint of model training, particularly for large-scale architectures, remains a concern, especially when models are trained without transparent reporting. Moreover, EO datasets often contain geographic, sensor-based, and socioeconomic imbalances, which can propagate biases into model predictions [17]. These biases risk producing results that are not representative of the true situation. Mitigation strategies include careful data curation, geographic rebalancing, and inclusive benchmarking practices.
Last but not least, we want to emphasize that while technical benchmarking is a necessary step, it is not a sufficient one. Public perceptions significantly influence policy decisions by shaping political priorities, public support, and the feasibility of implementation. Performance metrics alone do not guarantee real-world relevance or societal impact. While benchmarks offer a standardized way to measure methodological progress, they must be grounded in broader, mission-driven goals. As aligned with studies such as [35], the impactful deployment of geospatial FMs, especially in support of SDG-17, relies on collaboration with stakeholders, domain experts, and decision-makers who can translate model outputs into actionable insights for "Becoming Good at AI for Good". In practice, benchmarking and modeling represent just two components of a complete solution. Defining relevant problems, selecting appropriate metrics, involving domain experts, and clearly articulating the model’s purpose are equally important. FMs, unlike narrow models that suffer from “catastrophic forgetting”, offer unique advantages for long-term monitoring and adaptation to temporal shifts in the Earth system. Well-pretrained FMs enable lifelong learning and can generalize to new tasks with minimal further tuning.
Methods
Achieving the SDGs requires reliable tools to monitor progress across complex, interconnected systems. Despite the global availability of satellite data, assessing the real-world impact of development processes and their alignment with sustainability goals remains a major challenge [2]. We introduce SustainFM, a benchmark dataset for monitoring 16 SDGs using EO data. Table 4 provides the overview of SustainFM, where each entry corresponds to a distinct task aligned with a specific SDG-related application. To ensure broad applicability, SustainFM includes data from a variety of satellite sensors, spatial resolutions, geographies, and AI task types, including regression (Reg.), classification (Cls.), semantic segmentation (Seg.), and change detection (CD), capturing the diversity of real-world use cases. Given the broad spectrum of AI tasks associated with different SDGs, we attach task-specific decoders to each FM encoder. For semantic segmentation, we use a UperNet decoder; for change detection, we use a Siamese UperNet; and for classification and regression, we use a fully connected layer with global pooling. Optimization uses cross-entropy loss for classification, segmentation, change detection tasks, and mean squared error (MSE) loss for regression.
SustainFM is designed as a testbed for downstream tasks in FMs to enable systematic evaluation of performance, generalization, and energy efficiency across data modalities and learning tasks. SustainFM includes multiple EO sources, including Landsat, Sentinel-1/2, VIIRS, Gaofen-2, Google Earth, and PlanetScope, with spatial resolutions ranging from 0.5 m to 30 m. Geographically, the dataset spans globally, covering over 200 countries and regions, to minimize regional bias. The tasks tackle real-world challenges such as health, education, infrastructure, industrial activity, and disaster response. Below, we introduce the 16 building blocks of SustainFM, each aligned with a specific SDG, along with the model evaluation protocols.
Table 4: Description of the SustainFM Benchmark, including SDG-related tasks (applications) with sensor type, resolution, scale, AI task, patch size, and number of patches.
#SDG Application Sensor Resolution Scale AI Task Patch Size #Patches
[…]
2 Cropland Change Detection Gaofen-2 0.5-2m China CD $256\times 256256 × 256$ 2,400
[…]
4 Women Educational Attainment Regression Landsat+VIIRS 30m Africa Reg. $256\times 256256 × 256$ 117,062
[…]
6 Urban BOW Detection Gaofen-2 1m China Seg. $256\times 256256 × 256$ 1,645
[…]
8 Detection of Settlements without Electricity Sent-1+Sent-2+Landsat+VIIRS 10m Malawi Seg. $800\times 800800 × 800$ 98
[…]
10 Sanitation Index Regression Landsat+VIIRS 30m Africa Reg. $256\times 256256 × 256$ 89,271
[…]
12 Mining Change Detection Google Earth 1.2m Global CD $256\times 256256 × 256$ 71,711
[…]
14 Marine Pollution Detection Sent-2 10m Global Seg. $240\times 240240 × 240$ 2,803
[…]
16 War-induced Damage Classification Planet Scope 3m Regional Cls. $128\times 128128 × 128$ 13,460
SustainFM Benchmark Dataset
No Poverty (SDG-1): SustainFM includes an asset wealth index regression task to assess localised poverty using Demographic and Health Surveys (DHS) data from SustainBench [36]. Wealth is estimated at the cluster level using a principal component analysis (PCA) of household assets, yielding a scalar index averaged across surveyed households in each cluster. The dataset comprises 2,079,036 households across 86,936 clusters in 48 countries, based on DHS surveys conducted between 1996 and 2019.
Zero Hunger (SDG-2): SustainFM includes the Cropland Change Detection (CLCD) data set to support food security and sustainable agricultural land management [37]. CLCD comprises 600 paired satellite image samples (512×512 pixels) from Gaofen-2, captured in Guangdong Province, China, between 2017 and 2019. Each pair is annotated with binary labels that indicate cropland conversion, including changes to buildings, roads, lakes, and bare soil, key threats to agricultural sustainability.
Good Health and Wellbeing (SDG-3): SustainFM incorporates a children’s health regression task using DHS survey data from SustainBench [36]. Children’s health is measured via cluster-level child mortality rates, defined as the number of deaths per 1,000 children under age five within a defined spatial unit. The dataset comprises mortality rates for 105,582 clusters, derived from 1,936,904 children surveyed between 1996 and 2019.
Quality Education (SDG-4): SustainFM includes an analysis of women’s educational attainment, based on the DHS survey dataset from SustainBench [36]. The educational attainment metric is derived by calculating the cluster-level mean of education years for women aged 15-49. Education years are capped at 18 to mitigate outliers, a standard survey threshold. The dataset includes 122,435 cluster-level labels, computed from data on 3,013,286 women across 56 countries.
Gender Equality (SDG-5): SustainFM includes a women’s body mass index (BMI) regression task, using satellite imagery and survey data from SustainBench [36]. This dataset includes 94,866 cluster-level BMI labels, derived from 1,781,403 women of reproductive age (15–49), excluding pregnant individuals.
Clean Water (SDG-6): SustainFM includes an urban black and odorous water (BOW) detection task using the RSBD dataset [38], which comprises 329 multispectral images from the Gaofen-2 satellite in Yantai, China. This region has faced escalating pollution and freshwater shortages. Pixel-level annotations of BOW facilitate the training of segmentation models to detect contaminated water and support water quality management in urban settings.
Affordable and Clean Energy (SDG-7): SustainFM includes a solar farm identification task based on the GloSoFarID dataset [39], which provides 13,703 multispectral Sentinel-2 image samples with pixel-level annotations. The dataset includes global solar farm samples from 2017–2018 and U.S. samples from 2022–2023, covering installations at 10m spatial resolution. These EO-derived maps support the analysis of solar infrastructure expansion and guide decision-making in clean energy planning.
Decent Work and Economic Growth (SDG-8): SustainFM includes a task for detecting settlements without electricity using the DSE dataset [40], developed for the 2021 IEEE GRSS Data Fusion Contest. The dataset spans Malawi’s Dezda and Salima districts, covering 98 EO tiles of 800×800 pixels each ( 64 km²). It combines Sentinel-1, Sentinel-2, Landsat 8, and VIIRS nighttime imagery, labeling areas based on both residential presence and electricity access across four categorical classes. This task supports monitoring of energy poverty and associated economic disparities.
Industry, Innovation, and Infrastructure (SDG-9): SustainFM includes a smoke plume detection task based on the SmokePlume dataset [41], which supports model development for industrial impact assessment and environmental monitoring. The dataset builds on 624 industrial sites listed in the European Pollutant Release and Transfer Register, with Sentinel-2 Level-2A imagery from 2019. It includes 21,350 multispectral image patches (120 × 120 pixels; 1.2 × 1.2 km²), of which 3,750 contain visible smoke plumes, accompanied by pixel-level masks.
Reduced Inequalities (SDG-10): SustainFM includes an analysis of the sanitation index (unequal access to sanitary services), based on the DHS survey dataset from SustainBench [36]. The dataset utilizes a cluster-level toilet index to represent the sanitation index, ranked on a 1-5 scale, where 5 is the “highest quality”. The dataset comprises sanitation index for 89,271 clusters, computed from 2,143,329 households across 49 countries.
Sustainable Cities and Communities (SDG-11): SustainFM includes a building change detection task as a key proxy for urban dynamics by leveraging the LEVIRCD dataset [42]. The dataset captures significant urban development through Google Earth high-resolution imagery (0.5 m per pixel) across Texas, USA, from 2002 to 2018. It includes 637 image pairs (1024 × 1024 pixels), annotated with building additions and removals, encompassing diverse structures such as residential villas, high-rise apartments, garages, and warehouses.
Responsible Consumption and Production (SDG-12) SustainFM addresses environmental degradation and conserving natural resources through monitoring changes caused by mining activities globally based on the MineNetCD dataset [43]. MineNetCD includes over 70,000 bitemporal image patches (256 × 256 pixels) from 100 global mining sites across six continents, collected via Google Earth Engine (GEE). Pixel-level change masks annotate mining expansion to support impact assessments and the pursuit of responsible production.
Climate Action (SDG-13): The SustainFM benchmark includes a wildfire impact assessment task using the HLS Burn Scar Scenes dataset [44]. Based on the U.S. Monitoring Trends in Burn Severity (MTBS) database, this dataset combines burn scar shapefiles with Harmonized Landsat and Sentinel-2 (HLS) imagery acquired one to three months post-fire. It consists of 805 scenes, each measuring $512\times 512512 × 512$ pixels, with corresponding pixel-wise annotations of burn scars to support precise wildfire detection and post-event analysis.
Life below Water (SDG-14): SustainFM includes a task for detection of marine pollution—from oil slicks to plastic gyres—shedding light on eutrophication hotspots and floating debris through the Marine Debris and Oil Spills (MADOS) dataset [45]. This dataset comprises 174 Sentinel-2 scenes from 2015 to 2022, annotated with over 1.5 million pixels across diverse global waters and weather conditions. It captures pollution signatures at scale, spanning  47,000 m² of debris and  23 million m² of oil spills across 2803 patches ($240\times 240240 × 240$ px each, 2.4 km²).
Life on Land (SDG-15): SustainFM includes a flood detection task to advance disaster response and land management built on the OMBRIA dataset [46]. This dataset integrates bitemporal, multimodal satellite data (Sentinel-1 SAR and Sentinel-2 multispectral imagery) paired with expert-labeled ground truth from ESA’s Copernicus Emergency Management Service. Covering 23 major flood events worldwide (2017–2021), OMBRIA comprises 3,376 annotated images ($256\times 256256 × 256$ px).
Peace, Justice, and Strong Institutions (SDG-16): SustainFM contains a task to detect war-related damage using the GAZADeepDAV dataset [47], which features eight-band PlanetScope images at 3 m resolution. It comprises four pre- and four post-conflict scenes from Gaza in 2023, yielding 13,460 annotated patches ($256\times 256256 × 256$ px)—7,264 labeled as undamaged and 6,196 as damaged. GAZADeepDAV offers a unique foundation for developing EO tools that bolster transparency and resilience in conflict zones.
Evaluating Foundation Models on SustainFM Dataset
Preprocessing and Data Preparation
The SustainFM benchmark involves a variety of datasets with diverse spectral features, resolutions, and image sizes, which are designed for different AI applications, including segmentation, regression, and classification. This diversity brings difficulties to the evaluation of FMs on the SustainFM benchmark. Therefore, we preprocessed the dataset into a unified format to facilitate the data loading procedure to fine-tune and test the FMs. We resampled all the bands according to the highest resolution for the multispectral data, with different resolutions across the spectral bands. To ensure optimal performance and a convenient computation of the AI algorithms, we crop the satellite images of large sizes into small patches of around $256\times 256256 × 256$. To ensure data transparency and reproducibility, all preprocessed datasets have been uploaded to the Huggingface platform, which can be downloaded and reused in our implementation framework.
Foundation Models
Recent literature has witnessed rapid progress in developing FMs tailored to EO applications. We adopt the selection criteria proposed by PANGAEA [29] to ensure a rigorous and meaningful comparison, emphasizing reproducibility, methodological innovation, and scientific impact. Based on these criteria, we include 10 state-of-the-art FMs in our evaluation. Table 5 provides a comprehensive overview of the selected models, detailing their architectural scale, the datasets used for pretraining, and the associated EO sensors. These models represent diverse pretraining paradigms commonly employed in geospatial foundation modeling. Specifically, the selected FMs span three major pretraining strategies. Contrastive learning leverages auxiliary data in an unsupervised manner, aligning representations across different modalities to learn robust and transferable features. Masked image modeling employs a self-supervised approach, where portions of the input (i.e., patches of an image) are masked and the model is trained to reconstruct the missing content, thereby learning semantically meaningful representations. Supervised learning, by contrast, relies on labeled datasets to train the model directly for specific vision tasks, enabling the model to develop task-relevant perceptual capabilities. Below, we provide a detailed account of the ten FMs selected for evaluation, grouped according to their pretraining strategies: contrastive learning, masked image modeling, and supervised learning. These categories reflect the dominant paradigms currently shaping the development of general-purpose models in EO. An overview of the selected FMs is presented in Table 5, including their architectural backbones, pretraining datasets, learning strategies, model scales, and publication years.
  • Contrastive Learning: CROMA [18] aligns geographically and temporally co-registered SAR-optical image pairs using a contrastive objective, while also incorporating unimodal MIM losses to enhance modality-specific representations. RemoteCLIP [19] bridges the gap between vision and language by aligning satellite imagery with textual descriptions, resulting in visual representations enriched with high-level semantic concepts. SoftCon [20] introduces a multi-label supervision framework to learn cross-scene soft similarities, addressing challenges associated with rigid positive sample definitions and diverse semantic content in EO imagery.
  • Masked Image Modeling: DOFA [21] employs dynamic spectral adaptation to pretrain on multimodal EO data, explicitly incorporating physical characteristics of spectral bands based on wavelength information. GFM-Swin [22] applies cross-sensor reconstruction to learn generalizable representations across EO modalities. Prithvi [23] is the first FM pretrained on Harmonized Landsat and Sentinel-2 (HLS) data from NASA, integrating temporal dynamics during pretraining to capture seasonality and land cover evolution. SpectralGPT [10] adapts MIM for multispectral imagery, incorporating domain-specific mechanisms to accommodate spectral data’s unique structure and characteristics. ScaleMAE [24] addresses resolution heterogeneity in satellite data, learning robust representations from multimodal inputs at varying spatial scales. SSL4EO-S12 [25] serves as a foundational baseline FM, applying standard MIM techniques to Sentinel-1 and Sentinel-2 data in one of the earliest large-scale geospatial FM pretraining efforts.
  • Supervised Learning: SatlasNet [26] constructs a comprehensive multimodal dataset with diverse pretraining tasks, leveraging multi-task supervised learning to build an FM capable of capturing task-generalizable knowledge.
Table 5: Foundation models used in benchmarking, with architecture, pretraining data, strategy, size, and release year.
Model Arch. Pretrained EO Data Learning Strategy Params (M) Year
[…]
DOFA ViT DOFA [21] MIM 178.20 2024
[…]
Prithvi ViT Prithvi-HLS [23] MIM 153.28 2023
[…]
SatlasNet Swin-T SatlasPretrain [26] Supervised 128.57 2023
[…]
SpectralGPT ViT fMoW-S2 [48], BigEarthNet [49] MIM 614.75 2024
[…]
SoftCon ViT SSL4EO-S12 [25] Contrastive 242.19 2024
  • Acknowledgment: This study is supported by the 3D-ABC project funded by the Helmholtz Initiative Foundation Model Initiative. X.Z. discloses support for the research of this work from the National Natural Science Foundation of China under Grant No. 42371374. The authors would like to thank Caleb Robinson from Microsoft AI for Good and Prof. Guido Grosse from the Alfred Wegener Institute (AWI) for their valuable feedback, as well as Bo Li, Xuran Xiong, and Yuhang Li from Wuhan University of Science and Technology for their dedicated efforts in conducting the experiments.
  • Data availability: The data used for this study are sourced from publicly available databases, and detailed information about each variable’s source can be found in the ‘SustainFM Benchmark’ section of the Methodology. Furthermore, we provided the redistribution of the SustainFM benchmark based on the Huggingface platform: https://huggingface.co/HZDR-FWGEL.
  • Code availability: The code used for this study is built based on publicly available code snippets, and the full code framework is open-sourced on the Github platform: https://github.com/AI4RS/SustainFM.
  • Author contribution: Conceptualization: P.G. and W.Y. Methodology: P.G., W.Y., and X.Z. Coding and analysis: all authors. Writing-original draft: P.G. and W.Y. Writing-review and editing: all authors.
Supplementary Material
Figure S1: Model-wise comparison of mean F1-score across five datasets associated with different SDGs. This visualization highlights the performance differences among evaluated models, and provides insights into their generalization capabilities and robustness.
(a) Training Energy (kWh)
(b) Inference Energy (kWh)
(c) CO2 Emission (kg)
(d) GAZAdeepDAV Dataset
(e) GloSoFarID Dataset
(f) Ombria Dataset
(g) HLS Burns Dataset
(h) CLCD Dataset
Figure S2: Comprehensive evaluation of energy efficiency, carbon footprint, and predictive performance across diverse models and datasets. (a) Training energy consumption in kWh, (b) inference energy consumption in kWh, and (c) associated CO2 emissions in kg across all evaluated models. (d–h) Joint comparison of mF1 and CO2 emissions for five representative datasets (GAZAdeepDAV, GloSoForID, Ombria, HLS Burns, and CLCD), illustrating the trade-off between model accuracy and environmental sustainability.
(a) Effect of training data reduction on mF1 scores compared to full training.
(b) Differences in mF1 scores between full (100%) and reduced training samples.
Figure S3: Impact of training data reduction on model performance for HLS Burns. (a) mF1 scores achieved by each model when trained on progressively reduced subsets of the training data, compared to full training (100%). (b) Absolute differences in mF1 scores between full training and reduced-sample settings, highlighting model robustness to limited labeled data.
All experiments were conducted on high-performance computing clusters based on the PyTorch platform, utilizing an NVIDIA GeForce RTX 4090 graphics card with graphics memory of 24 GB.